The demand curve for Blivets (shown below) has the formula

 P =h(q) = 12-4

Where q is measured in thousand of Blivets and p is measured in dollars. We

Get Total revenue(measured in thousands of dollars) from this curve by the recipe

TR(q)=q*h(q). The total cost (measured in thousands of dollars ) of manufacturing Blivets

Is given by the formula TC(q)= q+1. Due to the limited supply of raw material, we

Cannot manufacture more than 3500 Blivets ( q=3.5)

 

A) write the formula for TR(q) and TR'(q) ( simplify your answer as much as possible)

b) find the value of q in the interval between q=0 and q=3.5 thousand blivets at which profit is greatest

c) find the formula for Total profit p(q) of selling q thousand blivets

d) b) find the value of q in the interval between q=0 and q=3.5 thousand blivets at which TR(q) reaches its largest value



1
Expert's answer
2021-05-10T15:24:10-0400

Given;

"p=12-4q"

"TC=q+1"



(a)

"TR=p\times q"

"=(12-4q)q"

"=12q-4q^2"


"MR=TR'"

"=\frac{d}{dq}(TR)"

"=\frac{d}{dq}(12q-4q^2)"

"=12-8q"


(b)

"profit (P)=TR-TC"

"=12q-4q^2-(q+1)"

"=-4q^2+11q-1"

Now,

"\frac{dP}{dq}=0" [first order condition for maximization of P)

"\frac{dP}{dq}(-4q^2+11q-1)=0"

"-8q+11=0"

"8q=11"

"q=\frac{11}{8}=1.375"

Thus, at the level of q=1.375(in thousand bilvets) maximum profit is realized


(c)

formula for total profit is

"P=TR-TC"

"=p\times q-TC"

"=(12-4q)q-(q+1)"

"=12q-4q^2-q-1"

"=11q-4q^2-1"

"=-4q^2+11q-1"


(d)

"TR=price\times quantity"

"=p\times(12-4q)"

"=12q-4q^2"

"\frac{dTR}{dq}=12-8q=0"

"q=\frac{12}{8}"

=1.5

Thus, at the level of q=1.5(in thousand bilvets) the TR reaches its maximum value.